Let Pk = {p(x)|p(x) is a polynomial of 

degree ≤ k with real coefficients}, for k €N.

Apply the Fundamental Theorem of 

Homomorphism to prove that p5/p3 is isomorphic to p1.


1
Expert's answer
2021-07-14T14:57:11-0400

Here Pk={p(x):f(x) is polynomial of degree "\le" k with real coefficient}

Let

f=p5"\to" P1 with the function defined by f(ax5+bx4+cx3+dx2+ex+1)=ax+b

for all ax5+bx4+cx3+dx2+ex+f "\in" P

where a,b,c,d,e,f are arbitary constants.

then it is clear that, f is homomorphism.

Because,

let p(x) (a1x5+b1x4+c1x3+d1x2+e1x+f1)"\in" P5

q(x)=(a2x5+b2x4+c2x3+d2x2+e2x+f2)"\in" P5

now,F(f(x)+q(x))=

f{(a1+a2)x5+(b1+b2)x4+(c1+c2)x3+(d1+d2)x2+(e1+e2)x+f1+f2}

=(a1+a2)x+(b1+b2)

=(a1x+b1)+(a2x+b2)

=f(p(x))+f(q(x)) "\forall" p(x),q(x)"\in" P5

f is homomorphism.

Also,

kerf={P(x)+P5:f(p(x))=0}

f(p(x))=0

a1x+b=0 "\forall" x"\in" R

a1=0,b1=0

p(x)=c1x3+d1x2+e1x+f "\in" P3

ker f=p3

also,

since f is a onto mapping

by homomorphism fundamental therom

P5\kerf=P1

P5/P3=P1